If the probability of being born on a specific day of the week is the same for all days of the week, about how many of 100 babies will probably have a Friday birthday?
Answer by stanbon(75887)

If the probability of being born on a specific day of the week is the same for all days of the week, about how many of 100 babies will probably have a Friday birthday?
-------------------
P(born on Friday) = 1/7
--------
# of 100 expected born on Friday = (1/7)100 = 14.29
---------------
Cheers,
Stan H.
===============